# Faster Convergence of Stochastic Accelerated Gradient Descent under Interpolation

### v1

@@ 3 Convergence of Stochastic AGD
Choosing $m$ to be one step of SGD with step-size $\eta_{k}$ yields the familiar updates of the standard version of stochastic AGD (see Nesterov [26, Eq. 2.2.20]):
@@
Suppose $f$ is $L$-smooth and $\mu$-strongly convex with $\mu>0$, $\eta_{\text{min}}\leq\eta_{k}<1/\mu$ almost surely for every $k\in\mathbb{N}$, and the primal update $m(\eta_{k},y_{k},\nabla f(y_{k},z_{k}))$ satisfies the sufficient progress condition (Eq. 8). If $\gamma_{0}=\mu$, then generalized stochastic AGD has the following rate of convergence:
@@
Suppose $f$ is $L$-smooth and $\mu$-strongly convex with $\mu\geq 0$, $\eta_{\text{min}}\leq\eta_{k}<1/\mu$ almost surely for every $k\in\mathbb{N}$, and the primal update $m(\eta_{k},y_{k},\nabla f(y_{k},z_{k}))$ satisfies the sufficient progress condition (Eq. 8). If $\gamma_{0}\in(\mu,3/\eta_{\text{min}})$, then generalized stochastic AGD has the following rate of convergence:
@@ 3.1 Specializations
If $f$ is $L$-smooth and $\mu$-strongly convex with $\mu>0$, strong growth holds, and $\eta_{k}=1/\rho L$, then stochastic AGD with $\gamma_{0}=\mu$ converges as,
@@ Acknowledgments
We would like to thank Frederik Kunstner, Victor Sanchez-Portella, and Sharan Vaswani for many insightful discussions.
@@
[…]
Data Availability
This paper does not rely any data, public or private. Derivations for all theoretical results are included in the appendix.

### v2

@@ Abstract
This preprint has a significant bug in its proofs. In particular, the claim that the generalized stochastic AGD scheme in Eq. 5 with the map $m$ set to be stochastic gradient update is equivalent to the standard “momentum” version of stochastic AGD in Eq. 6 does not hold. Unfortunately, the bug invalidates the main conclusion of the paper — namely that the dependence on the strong growth constant can be improved from $\rho$ to $\sqrt{\rho}$ for stochastic Nesterov acceleration. We are currently do not know if this issue can be corrected to obtain the desired $\sqrt{\rho}$ dependence or if $\rho$ is in fact tight. We will continue to explore this issue and post an updated version of the preprint if we obtain any new results. Note that all results hold as stated for the semi-stochastic scheme in Eq. 5. Please see Appendix D for further details on the bug.
@@ 3 Convergence of Stochastic AGD
Choosing $m$ to be one step of SGD with step-size $\eta_{k}$ yields the familiar updates of the standard version of stochastic AGD (see Nesterov [26, Eq. 2.2.20]). (Warning: The preceding claim is not true. The use of a deterministic gradient in the update for $v_{k+1}$ in Eq. 5 breaks the standard argument that this scheme is equivalent to the “momentum” form of stochastic AGD in Eq. 6. Please see Appendix D for more details.)
@@
(Warning: This result only holds for the scheme in Eq. 5. It does not hold for stochastic AGD.) Suppose $f$ is $L$-smooth and $\mu$-strongly convex with $\mu>0$, $\eta_{\text{min}}\leq\eta_{k}<1/\mu$ almost surely for every $k\in\mathbb{N}$, and the primal update $m(\eta_{k},y_{k},\nabla f(y_{k},z_{k}))$ satisfies the sufficient progress condition (Eq. 8). If $\gamma_{0}=\mu$, then generalized stochastic AGD has the following rate of convergence:
@@
(Warning: This result only holds for the scheme in Eq. 5. It does not hold for stochastic AGD.) Suppose $f$ is $L$-smooth and $\mu$-strongly convex with $\mu\geq 0$, $\eta_{\text{min}}\leq\eta_{k}<1/\mu$ almost surely for every $k\in\mathbb{N}$, and the primal update $m(\eta_{k},y_{k},\nabla f(y_{k},z_{k}))$ satisfies the sufficient progress condition (Eq. 8). If $\gamma_{0}\in(\mu,3/\eta_{\text{min}})$, then generalized stochastic AGD has the following rate of convergence:
@@ 3.1 Specializations
(Warning: This result only holds for the scheme in Eq. 5. It does not hold for stochastic AGD.) If $f$ is $L$-smooth and $\mu$-strongly convex with $\mu>0$, strong growth holds, and $\eta_{k}=1/\rho L$, then stochastic AGD with $\gamma_{0}=\mu$ converges as,
@@ Acknowledgments
We would like to thank Frederik Kunstner, Victor Sanchez-Portella, and Sharan Vaswani for many insightful discussions. We thank Chia-Yu Hsu for their help in discovering the bug in this preprint.
@@ Appendix C Comparison to Existing Rates: Proofs
Appendix D Theoretical Issues in the Preprint
As stated in the preamble, the optimization schemes given in Eq. 5 and Eq. 6 are not equivalent as Section 3 claims. While the formal equivalence of the deterministic versions of these two algorithms is proved by Nesterov [26, Eq. 2.2.20], their argument fails in our setting because Eq. 5 combines a deterministic estimating sequence with stochastic gradient updates. In particular, the deterministic estimating sequence implies deterministic gradient updates for $v_{k}$, which cannot reconciled with the stochastic updates for $x_{k}$ when using the standard argument. See Section D.1 for sketch of how the proof fails.
This bug does not falsify all results in this preprint. Since the estimating sequence form of AGD in Eq. 5 is the target of our theoretical analysis in Section 3, all of the claimed convergence rates continue to hold for this scheme. They fail only for the momentum version of stochastic AGD in Eq. 6. However, the estimating sequence form of AGD is not truly stochastic because, as noted, Eq. 5 requires deterministic gradient updates for $v_{k}$. Thus, we are only able to prove much weaker results than the original claims in the preprint.
There are several potential ways to resolve this problem. One approach is to change the estimating sequence to use stochastic gradient updates. That is, change Eq. 9 to,
$\lambda_{k+1}$ $=(1-\alpha_{k})\lambda_{k}$ (33)
$\phi_{k+1}(w)$ $=(1-\alpha_{k})\phi_{k}(w)+\alpha_{k}(f(y_{k})+\left\langle\nabla f(y_{k},z_{k}),w-y_{k}\right\rangle+\frac{\mu}{2}\|w-y_{k}\|_{2}^{2}).$
Since $v_{k+1}$ is the minimizer of $\phi_{k+1}$, this results in a stochastic gradient update for $v_{k+1}$ in Eq. 5. Analyzing this fully stochastic version is straightforward if strong growth is used to control the $\|\nabla f(y_{k},z_{k})\|_{2}^{2}$ term which now appears in $\phi_{k+1}^{*}$. Disappointingly, this approach leads to the $\rho$ dependence (as opposed to $\sqrt{\rho}$) previously established by Vaswani et al. [22].
Another approach is to retain a deterministic estimating sequence while still modifying Eq. 5 to use a stochastic update for $v_{k}$. In that case, we obtain two different $v_{k}$ sequences: the true minimizers of $\phi_{k}$ and a sequence of unbiased stochastic estimates $\tilde{v}_{k}$ maintained by the algorithm. While promising, this approach leads to the expectation $\mathbb{E}\left[\left\langle\nabla f(y_{k}),v_{k}-\tilde{v}_{k}\right\rangle\right]$ in the analysis of Proposition 5. The issue here is that $\tilde{v}_{k}$ is correlated with $\nabla f(y_{k})$ as $y_{k}$ is computed from $\tilde{v}_{k}$, meaning this expectation does not resolve to zero. It is not clear whether or not a more careful analysis of this term can yield the desired $\sqrt{\rho}$ dependence.
Another option is that obtaining a $\sqrt{\rho}$ dependence for stochastic AGD is not possible. That is, there is a lower bound showing that the $O(\rho\sqrt{L/\mu}\log(1/\epsilon))$ complexity proved Vaswani et al. [22] is in fact tight. We are investigating this possibility.
D.1 Failed Equivalence Argument
Now we show how the standard equivalence argument fails. Our goal is to eliminate the $v_{k}$ sequence by writing $y_{k+1}$ as a function of $y_{k}$ and $w_{k+1}$. Recalling that,
we can re-arrange to obtain the following expression for $v_{k}$,
$v_{k}=\frac{1}{\alpha_{k}\gamma_{k}}((\gamma_{k}+\alpha_{k}\mu)y_{k}-\gamma_{k+1}w_{k}).$
Then, starting from the definition of $v_{k+1}$ in Eq. 5 and substituting in this value for $v_{k}$, we obtain,
$v_{k+1}$ $=\frac{1}{\gamma_{k+1}}\left[\frac{(1-\alpha_{k})}{\alpha_{k}}((\gamma_{k}+\alpha_{k}\mu)y_{k}-\gamma_{k+1}w_{k})+\alpha_{k}\mu y_{k}-\alpha_{k}\nabla f(y_{k})\right]$
$=\frac{(1-\alpha_{k})\gamma_{k}}{\alpha_{k}\gamma_{k+1}}y_{k}+\frac{\mu}{\gamma_{k+1}}y_{k}-\frac{(1-\alpha_{k})}{\alpha_{k}}w_{k}-\frac{\alpha_{k}}{\gamma_{k+1}}\nabla f(y_{k})$
$=\frac{1}{\alpha_{k}}y_{k}-\frac{(1-\alpha_{k})}{\alpha_{k}}w_{k}-\frac{\eta_{k}}{\alpha_{k}}\nabla f(y_{k})$
$=w_{k}+\frac{1}{\alpha_{k}}\left(y_{k}-\eta_{k}\nabla f(y_{k})-w_{k}\right)$
$=w_{k}+\frac{1}{\alpha_{k}}\left(w_{k+1}-w_{k}\right)+{\color[rgb]{1,0,0}\frac{\eta_{k}}{\alpha_{k}}\left(\nabla f(y_{k},z_{k})-\nabla f(y_{k})\right)}.$
The proof now proceeds by substituting this expression for $v_{k+1}$ into the equation for $y_{k+1}$. Although the error term $\frac{\eta_{k}}{\alpha_{k}}\left(\nabla f(y_{k},z_{k})-\nabla f(y_{k})\right)$ is expectation zero, it cannot be removed. This breaks the proof that the schemes in Eq. 5 and Eq. 6 are equivalent.